Can someone help me?
What is the algebriac equation for:
5 times a number minus ten is equal to the sum of six and three
Thanks, Lisa

Answer by checkley71(8403)

5X-10=6+3 OR 5X=10+6+3 OR 5X=19 OR X=19/5 OR 3 4/5 OR 3.8
PROOF 5*3.8-10=6+3 OR 19-10=9 OR 9=9